OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (radar):

From looking at the "z" terms in the last two equations, I note two important facts: Fact 1 they are opposite signs Fact 2 one is twice the value of the other (disregarding signs). Now to activate my fact analyzer: Now activated. If i multiply the second equation all terms by 2 I get the following: 4x-4y-2z=2 well all I need to do now is to add: 4x-4y-2z=2 2x-y +2z=-15 ------------------results in 6x-5y = -13 resulting equation. a